# Analysis of Failures and Risks in Deep Learning Model Converters: A Case Study in the ONNX Ecosystem

### v3

@@ 6. Theme 3: Investigating Deeper Causes
Does ONNX evolution affect converter failure rates? RQ3: Does ONNX evolution affect converter failure rates? We hypothesize H: Changes in ONNX operator sets are correlated with increased defects. This hypothesis is based in data from our survey and failure analysis. In the survey, 19% of respondents reported changing ONNX versions as a possible solution (section 4, Table 5). This suggests that conversion defects (not just compatibility issues) may be localized by version. In the failure analysis, we found that crashes were largely due to Incompatibilities or Type Problems (section 5.2, Table 10). These failure modes relate to API compatibility and conversion correctness, which can be affected by changes in the ONNX opset.
RQ4: Do model types affect converter failure rates? We hypothesize H: Failures are caused by model structure, i.e., models with particular layers (node) or layer sequences are more prone to defects. This hypothesis is primarily based on our failure analysis. The majority of failures occurred during Node Conversion (Table 8), indicating that nodes and node sequences may be problematic. Survey data also had a relevant anecdote: one respondent solved conversion issues by re-implementing models using pure tensor operations, implying that conversion may fail due to the presence of certain operations.
@@ 6.1. RQ3: Does ONNX evolution lead to failures?
6.1.1. Method
We test the hypothesis by checking whether larger changes in the ONNX specification correlate with greater incidence of failures in DL model converters. To obtain the size of changes to the ONNX specification, we count the number of operator changes per release in the ONNX release notes (Figure 5). We approximate the number of failures per release using the GitHub issue creation times of the failures sampled for RQ2. We then attribute each issue to the nearest previous ONNX release.
We note four ways in which this measurement approximates: (1) DL model converters lag behind ONNX releases (this might cause a failure to be mis-attributed to another reuse, i.e., offset in time) ; (2) Failures might be in any ONNX available release, not just the most recent (possibly inflating the failure rate of a given release); (3) our failure analysis data were randomly sampled, possibly under-sampling certain time windows (though we note that our sample comprises 30-50% of qualifying issues); (4) as noted in the user survey (section 4.2.3), users may simply revert to a previous operator set without opening a bug report. For items 1 and 2, issues do not reliably include ONNX versions, so no more accurate data is available. For item 3, we use the data sampled during RQ2 so that we can check incidences by cause (incompatibility and type problems).
We test the hypothesis qualitatively and quantitatively. Qualitatively, we inspect a visualization of the hypothesized trend. We also measure the relationship, assessing the correlation in the number of changes in an ONNX release and the number of failures between its release and the next. We use the Spearman correlation, which is a commonly-used and robust metric for measuring a monotonic relationship between two variables (Fenton and Bieman 2014).
@@ 6.2.1. Method for macro analysis
At the macro scale, we test DL converters using two types of models and then analyze the failure-inducing inputs. We used a differential testing approach (Figure 7). Differential tests need (1) inputs, and (2) a difference criterion (McKeeman 1998).
@@
Real Models: We used the HFTorrent dataset from Jiang et al. (Jiang et al. 2023d). At time of experiment, this was the largest available set of real-world DL models, containing 63,182 pre-trained models collected from the HuggingFace model registry. We filtered for the 1,761 models with both PyTorch and TensorFlow versions, allowing comparison between the two converters on similar inputs. These models represent 112 distinct full architectures (58 backbones).
Synthetic Models: We systematically generated synthetic DNNs for conversion, using the NNSmith tool (Liu et al. 2023). NNSmith generates random valid TensorFlow and PyTorch DNNs that use operators supported by ONNX. We added a parameter to NNSmith to generate DNNs of a given size (# nodes). We then systematically generated DNNs containing between 15-100 nodes in increments of 5. For each node count, we generated DNNs until two conditions were met: (1) All NNSmith-supported operators appeared at least once in the family; and (2) 100 distinct models were created.
@@
There are other DL model generation approaches, e.g., MUFFIN (Gu et al. 2022), COMET (Li et al. 2023), and LEMON (Wang et al. 2020). These tools generate inputs by mutating seed models, typically real models. We viewed HFTorrent as a sufficient source of real models, and synthesized with NNSmith as a complementary approach.
@@
In both cases, to measure model misbehavior we used a simple community-accepted approach: the distance measure and threshold used by the PyTorch exporter. This rule is that between the original and ONNX-converted models, in the inference result tensors, the maximum absolute element-wise difference should be (PyTorch 2023).
We considered alternatives to measure behavioral differences. A more general notion of tolerance or “acceptable error” could be used, but the choice of tolerance is not well established — the PyTorch verification tool uses , NNSmith uses (Liu et al. 2023). Openja et al. (Openja et al. 2022) proposed using model accuracy and robustness, but this method requires training each model on a suitable dataset. Training substantially increases the cost of the measurement, and the dataset requirement limits the types of models that can be generated.
@@ 6.2.2. Results for macro analysis
Crashes: Synthetic models exhibit difficulties when converting to ONNX and loading converted models to the ONNX Runtime. We find 802/3,544 (23%) Unsuccessful Conversions, almost entirely in synthetic models on tf2onnx. These crashes correspond to 4 unique crash locations. Of the 2,742/3,544 (77%) synthetic models that successfully convert, only 1,244/2,742 (45%) successfully load into ONNX Runtime. We observed 11 unique ONNX Runtime errors, of which 6 appear to correspond to open GitHub issues (vbogach 2022; josephrocca 2021; BowenBao 2023; rafaelagrc 2022; fatcat z 2022; 12sf12 2022). We disclosed the rest to the engineering teams. For most disclosures, we have not yet received a response. For one disclosure, which described a model that causes the libc++ abi to terminate unexpectedly during inference with ONNX Runtime, the response was that it had been (unintentionally) fixed in the development branch. It still affects older releases.
@@ 6.2.3. Method for micro analysis
For operator types, we measure the operators present in each converted model, out of the set of available operators in the ONNX opset. For operator sequences, we extract the simple paths through each model. To illustrate, the ONNX model in Figure 2 shows three operators (ReduceMax_0, ArgMax, and ReduceMax) and two simple linear paths (ReduceMax_0ArgMax and ReduceMax_0 ReduceMax). Models with similar architectures may have simple paths that are largely identical, inflating the number of shared sequences. For example, the sequences aaaabd and aaaabc are identical except for the last element. To deal with cases we further reduce the common sequences to the smallest shared subsequences, i.e., we recursively find the longest common subsequences for our sequences until we cannot find any smaller subsequences (Cormen et al. 2022). After reduction we find that sequences are between 3 to 5 operators long.
@@ 7.1. Validating DL Model Converters
Converter test suites share little in common with failing models, often missing critical model structures (section 6.2.4). Specifically, model converters need better testing techniques for behavioral changes. Though this critical failure mode occurred in 1/3 of historical failures (section 5.2) and affected 6% of the models we tested (section 6.2.2), the existing test suite models share little in common with models affected by this failure mode. For example, only 37 operator sequences are shared between mismatching models and test suite models (Table 12). Introducing operator sequence coverage metrics may help improve the quality of test suites, to promote the diversity of the models tested. These metrics could also be incorporated into fuzzers.
Identifying behavioral changes after model conversion is critical, this is typically done with an end-to-end test. End-to-end testing carries with it the need to assess the outcome: is the converted model acceptable, and how much error is tolerable (Jiang et al. 2023b; Face 2023)? Engineers currently use a variety of heuristic tolerances derived from experience (section 6.2.1). Theoretical and empirical examination of the expected and acceptable tolerance would benefit DL model converter testing.
@@ 7.2.1. DL Model Converters
Openja et al. evaluated DL model converters by converting 5 models (Openja et al. 2022). We extend their findings in two directions. First, we conducted a failure analysis (section 5) to study how converters fail. Second, we convert models systematically (section 6) and analyze them to understand why they fail. We considered both realistic and synthetic models. For realistic models, rather than picking 5 real-world models, we used 1,605 models from a large model corpus. For synthetic models, we adapted a DL model generation tool to conduct a bounded systematic exploration of converter behaviors. We omitted measurements of model size, adversarial robustness, and prediction accuracy, to focus instead on measuring the common failure modes (crashing and behavioral differences) identified in our failure analysis. Our analysis reveals that converters can successfully convert many real model but synthetic models are more prone to failure, moreover analyzing the synthetic models reveals that particular operator sequences co-occur with failures.
@@ 7.2.2. DL Compilers
Shen et al. studied failures in DL compilers (Shen et al. 2021a), including the components that load models from DL frameworks and from common intermediary formats such as ONNX. Many of their results generalized to DL model converters: (1) we were able to adapt their taxonomies of failure symptoms and causes to DL model converters (section 5.1.1, Table 7); and (2) as shown in Table 9, we found a similar distribution of failure symptoms for DL model converters. However, the causes of failures between the two contexts differed: as noted in section 5.2.3, Incompatibilities were more common in DL model converters, while Algorithmic Errors were more common in DL compilers.
@@ 7.2.3. Methodological Innovations on the “Bug Study”
Our study began with a survey that offered valuable insights into the problem domain, guiding our design of the subsequent failure study, including the selection of the interoperability tool and the creation of a failure taxonomy (section 5). This methodological progression, from survey to failure study, and then to hypothesis testing about the failures, diverges from the typical failure study in software engineering. Those studies typically focus on the “middle step” (Amusuo et al. 2022a), understanding the root causes of defects and the conditions under which they manifest (e.g., (Islam et al. 2019b; Zhang et al. 2018, 2019)). Our hypothesis testing phase, in particular, adds understanding by providing an estimate of failure rates, an aspect seldom addressed in standard failure studies. We suggest that this holistic approach, integrating a survey, failure study, and hypothesis testing, may be a useful methodology for future failure studies.
@@ 8. Threats to Validity
Construct: We mitigated definitional concerns for RQ2 by adapting existing taxonomies in our failure analysis. We defined failures consistently with other works (GitHub issues closed with a repair), though we note there are other means of failure discovery (Aranda and Venolia 2009). To answer RQ3, we conservatively defined failure in terms of clear misbehaviors — when ONNX converter output is incompatible with ONNX Runtime, or the model’s behavior changes. This may mask some “Crash” failures in the ONNX converter. In addition, different measures of behavioral difference are possible, such as the L1- or L2- distance. We used the measure recommended by PyTorch.
Internal: Our failure characterization in RQ2 was manual. We mitigated subjectivity via interrater agreement, on a pilot sample and a subsequent tranche during the full analysis (section 5.1.2). In section 6.1.1 we noted several approximations in our test of hypothesis H. We tested hypothesis H indirectly, using a frequency analysis rather than directly testing the discovered subsequences.
External: For RQ1, the surveyed population (HuggingFace users) may not fully reflect all ONNX users. However, it is an appropriate population, and we sampled at a confidence level of 90%. For RQ2-4, we examined two DL model converters for one interoperability framework (ONNX). Our results may not extend to other model converters nor other interoperability tasks. As mitigation, our data suggest that ONNX is the most popular DL interoperability tool. For ONNX, our results were similar across the two converters, suggested generalizability in two dimensions: DNN modeling approaches, and converter owners (section 5.1.1). With regard to our methodology for RQ3-4, we generated synthetic models with the NNSmith tool. Other methods (Gu et al. 2022; Li et al. 2023; Wang et al. 2020) might have different results.
@@ 9. Conclusion
DL model converters play a crucial role in providing interoperability between DL frameworks, registries, and runtimes. Understanding the nature of failures in DL model converters enables engineers to make informed decisions when relying on them. We conducted the first failure analysis of DL model converters, considering the PyTorch and TensorFlow converters for the popular ONNX intermediate representation. The most common symptoms of failure are crashes and, perhaps more concerningly, models that misbehave on certain inputs. Of the five stages of a typical model converter, one stage — node conversion — accounts for 75% of the defects. A deductive description of the causes of erroneous converter behavior remains elusive — individual operators are not predictive of failure; sequences of operators may be correlated. Our findings suggest that in ONNX, engineers can rely on model converters but should validate the result for behavioral consistency. Through a mix of positive and negative results, we exposed several directions for further improvement of DL model converters. The main opportunities are new measurements: a new architectural coverage measure for a DL model converter, and a refined measure of tolerance after conversion.
10. Acknowledgements
This work is supported by Cisco, Google, and by NSF #2107020, #2107230, #1813935, and #2104319.
Appendix A Summary of Appendices
Appendix A: Appendix Summary.
Appendix B: Taxonomies and Extended Tables from section 5.
Appendix C: Additional Tables from section 6.
Appendix B Theme 2 – Failure Analysis (Of ONNX)
B.1. Taxonomies
| Symptom | Definition |
| Crash | The model converter terminates unexpectedly during conversion, usually producing an error message. |
| Wrong Model | The model converter behaves unexpectedly without crashing, producing a wrong intermediate or final result. |
| Bad Performance | The time cost or memory consumption of the model converter is much larger than developer/user expectations (e.g., the performance achieved on the previous version during regression testing or performance achieved by certain hardware). |
| Build Failure | Installation of model converter or dependencies fails. |
| Unreported | The symptom cannot be identified by analyzing the corresponding issues including code changes, discussions, and related issues. |
| Hang | This symptom means that the converter keeps running for a long period of time without producing the expected result. |
B.2. Symptom and Cause Data
Appendix C Theme 3 – Deeper Causes
Table 15 contains results of model operator analysis as described in section 6.2.3.
| Causes | TF | PT | Total | Comp. to DL compilers (Shen et al. 2021a) | |
| Incompatibility | External | 23 | 32 | 55 (28%) | 16 (4%) |
| Internal | 2 | 0 | 2 (1%) | 5 (1%) | |
| Resource | 0 | 0 | 0 (0%) | 4 (1%) | |
| Type Problem | Node | 21 | 25 | 46 (23%) | 13 (4%) |
| Conventional | 3 | 2 | 5 (3%) | 20 (6%) | |
| Tensor | 1 | 2 | 3 (2%) | 42 (12%) | |
| Algorithmic Error | 18 | 6 | 24 (12%) | 59 (16%) | |
| Shape Problem | 9 | 12 | 21 (11%) | 67 (19%) | |
| API Misuse | 6 | 6 | 12 (6%) | 35 (10%) | |
| Others | 17 | 15 | 32 (16%) | 98 (27%) | |
| Total | 100 | 100 | 200 (100%) | 359 (100%) | |
| Model Type | tf2onnx | torch.onnx |
| Mismatched | 54 | 58 |
| Correct | 52 | 59 |
| Test Suite | 54 | 16 |
| 0 | 1 | |
| 2 | 0 | |
| 21 | 46 | |
| 21 | 4 |
| Crash | Wrong Model | Bad Performance | Build Failure | Unreported | Total | |||||||
| PT | TF | PT | TF | PT | TF | PT | TF | PT | TF | PT | TF | |
| API Misuse | 5 | 6 | 1 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 6 | 6 |
| Incompatibility | 28 | 19 | 3 | 4 | 0 | 0 | 0 | 0 | 1 | 2 | 32 | 25 |
| Type Problem | 14 | 8 | 13 | 17 | 1 | 0 | 0 | 0 | 1 | 0 | 29 | 25 |
| Shape Problem | 4 | 5 | 7 | 4 | 0 | 0 | 0 | 0 | 1 | 0 | 12 | 9 |
| Incorrect Numerical Computation | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 |
| Incorrect Assignment | 2 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 2 | 0 |
| Incorrect Exception Handling | 1 | 1 | 1 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 2 | 1 |
| Misconfiguration | 0 | 1 | 0 | 0 | 0 | 0 | 2 | 2 | 0 | 1 | 2 | 4 |
| Algorithmic Error | 3 | 4 | 3 | 10 | 0 | 2 | 0 | 0 | 0 | 2 | 6 | 18 |
| Typo | 2 | 0 | 0 | 0 | 0 | 0 | 0 | 1 | 0 | 0 | 2 | 1 |
| Testing | 1 | 0 | 1 | 0 | 0 | 0 | 0 | 0 | 1 | 1 | 3 | 1 |
| Others | 2 | 6 | 1 | 0 | 0 | 0 | 0 | 0 | 1 | 4 | 4 | 10 |
| Total | 62 | 50 | 30 | 35 | 1 | 2 | 2 | 3 | 5 | 10 | 100 | 100 |
| Causes | Definition |
| API Misuse | High-level category that encapsulates Wrong API, Condition missing/redundancy, and API missing/redundancy. |
| Wrong API | The developer or user uses the wrong API or wrong arguments in an API. |
| Condition missing/redundancy | Developer fails to use (or redundantly uses) a condition check for an API. |
| API missing/redundancy | Developers fail to use (or redundantly use) an API. |
| Incompatibility | High-level category that encapsulates Internal, External, and Resource. |
| Internal | There are API compatibility issues within a model converter caused by API evolution. |
| External | There are API compatibility issues between a model converter and third-party libraries (e.g., ONNX, TensorFlow, ONNX Runtime). |
| Resource | There are compatibility issues between model converters and external resources. The characteristics of the target device are incompatible with the model converter. |
| Type Problem | |
| Node | A model converter works on a computational graph, where nodes represent the atomic DL operators (such as convolution and pooling) and edges represent the tensors. Each node takes zero or more tensors as input and produces a tensor as output. |
| Tensor | This relates to the types of tensors. Specifically, a tensor is a multi-dimensional matrix containing elements of a single data type. |
| Conventional | There are also conventional variables widely used in the development of traditional software systems. This subcategory refers to the problem involving the types of conventional variables. |
| Shape Problem | This cause covers, broadly, the shape of inputs and outputs a node can have. These occur during the operation of shape matching, shape transformation, shape inference, layout transformation, etc. |
| Tensor Shape Problem | This is related to tensor shape or layout. The tensor shape is the number of elements in each dimension. Layout describes how the shape is represented in memory. |
| Incorrect Numerical Computation | This involves incorrect numerical computations, values, or usage, such as incorrect operators or operands, dividing by 0, and missing or redundant operands. |
| Incorrect Assignment | This involves a variable being incorrectly initialized or assigned, or a variable lack initialization. |
| Incorrect Exception Handling | This category occurs due to incorrect exception handling. For example, an exception is not thrown when it should be, an example is thrown when it should be, or an incorrect/imprecise exception message is thrown. |
| Misconfiguration | This category occurs due to misconfiguration in the model converter. This includes dependency versioning and misconfigured environments. |
| Concurrency | This is a result of incorrect operations on concurrency-oriented structures (e.g., locks, threads, and critical regions). |
| Algorithmic Error | High-level category that encapsulates Incorrect Optimization and Incorrect Tracing. |
| Incorrect Optimization | There is an issue with optimizations that occur in the model converter. Examples: incorrect fusing, dead-code elimination, et cetera. |
| Incorrect Tracing | There is an issue with the tracing of DL models. This occurs in frameworks that use dynamic models. Dynamic models need to be “traced” to find the nodes in the computational graph prior to the conversion. |
| Typo | This is due to trivial mistakes (e.g. slips (Norman 2013)) by developers, e.g., “default” is mistakenly written as “defualt”. |
| Testing | This class contains issues that are a result of incorrect, outdated, or missing documentation. |
| Documentation | This class contains issues related to software tests such as unit tests. This covers “flaky” tests, missing tests, broken tests, or the addition of new tests. |
| Others | This cause is relegated to causes that occur infrequently and do not belong to the other causes. |

### v4

@@ 6. Theme 3: Investigating Deeper Causes
RQ3: Does ONNX evolution affect converter failure rates? We hypothesize H: Changes in ONNX operator sets are correlated with increased defects. This hypothesis is based in data from our survey and failure analysis. In the survey, 19% of respondents reported changing ONNX versions as a possible solution (Table 5). This suggests that conversion defects (not just compatibility issues) may be localized by version. In the failure analysis, we found that crashes were largely due to Incompatibilities or Type Problems (Table 10). These failure modes relate to API compatibility and conversion correctness, which can be affected by changes in the ONNX opset.
RQ4: Do model types affect converter failure rates? We hypothesize H: Failures are caused by model structure, i.e., models with particular layers (node) or layer sequences are more prone to defects. This hypothesis is primarily based on our failure analysis. The majority of failures occurred during Node Conversion (Table 8), indicating that nodes and node sequences may be problematic. Survey data also had a relevant anecdote: one respondent solved conversion issues by re-implementing models in pure tensor operations, implying that conversion may fail due to the use of certain operations.
@@ 6.1. RQ3: Does ONNX Evolution Lead to Failures?
6.1.1. Methodology
We test the hypothesis by checking whether larger changes in the ONNX specification correlate with greater incidence of failures in DL model converters. To obtain the size of changes to the ONNX specification, we count the number of operator changes (additions or updates) per release in the ONNX release notes (Figure 5). We approximate the number of failures per release using the GitHub issue creation times of the failures sampled for RQ2. We then attribute each issue to the nearest previous ONNX release.
We note four ways in which this measurement approximates: (1) DL model converters lag behind ONNX releases (this might cause a failure to be mis-attributed to another release, i.e., offset in time); (2) Failures might be in any ONNX available release, not just the most recent (possibly inflating the failure rate of a given release); (3) our failure analysis data were randomly sampled, possibly under-sampling certain time windows (though we note that our sample comprises 30-50% of qualifying issues); (4) as noted in the user survey (section 4.2.3), users may simply revert to a previous operator set without opening a bug report. For items 1 and 2, issues do not reliably include ONNX versions, so no more accurate data is available. For item 3, we use the data sampled during RQ2 so that we can check incidences by cause (incompatibility and type problems).
We test the hypothesis qualitatively and quantitatively. Qualitatively, we inspect a visualization of the hypothesized trend. We also measure the relationship, assessing the correlation in the number of changes in an ONNX release and the number of failures between its release and the next. We use the Spearman correlation, which is a commonly-used and robust metric for measuring a monotonic relationship between two variables (Fenton and Bieman, 2014).
@@ 6.2.1. Method for Macro Analysis
At the macro scale, we test DL converters using two types of models and then analyze the failure-inducing inputs. We used a differential testing approach (Figure 7). Differential tests need (1) inputs, and (2) a difference criterion (McKeeman, 1998).
@@
Real Models: We used the HFTorrent dataset from Jiang et al. (Jiang et al., 2023). At time of experiment, this was the largest available set of real-world DL models, containing 63,182 pre-trained models collected from the HuggingFace model registry.111Future measurements could use the PeaTMOSS dataset (Jiang et al., 2024b) or its successors. We filtered for the 1,761 models with both PyTorch and TensorFlow versions, allowing comparison between the two converters on similar inputs. These models represent 112 distinct full architectures (58 backbones).
Synthetic Models: We systematically generated synthetic DNNs for conversion, using the NNSmith tool (Liu et al., 2023). NNSmith generates random valid TensorFlow and PyTorch DNNs that use operators supported by ONNX. We added a parameter to NNSmith to generate DNNs of a given size (# nodes). We then systematically generated DNNs containing between 15-100 nodes in increments of 5. For each node count, we generated DNNs until two conditions were met: (1) All NNSmith-supported operators appeared at least once in the family; and (2) 100 distinct models were created.
@@
There are other DL model generation approaches, e.g., MUFFIN (Gu et al., 2022), COMET (Li et al., 2023), and LEMON (Wang et al., 2020). These tools generate inputs by mutating seed models, typically real models. We viewed HFTorrent as a sufficient source of real models, and synthesized with NNSmith as a complementary approach.
@@
In both cases, to measure model misbehavior we used a simple community-accepted approach: the distance measure and threshold used by the PyTorch exporter. This rule is that between the original and ONNX-converted models, in the inference result tensors, the maximum absolute element-wise difference should be (PyTorch, 2023).
We considered alternatives to measure behavioral differences. A more general notion of tolerance or “acceptable error” could be used, but the choice of tolerance is not well established — the PyTorch verification tool uses , NNSmith uses (Liu et al., 2023). Openja et al. (Openja et al., 2022) proposed using model accuracy and robustness, but this method requires training each model on a suitable dataset. Training substantially increases the cost of the measurement, and the dataset requirement limits the types of models that can be generated.
@@ 6.2.2. Results for Macro Analysis
Crashes: Synthetic models exhibit difficulties when converting to ONNX and loading converted models to the ONNX Runtime. We find 802/3,544 (23%) Unsuccessful Conversions, almost entirely in synthetic models on tf2onnx. These crashes correspond to 4 unique crash locations. Of the 2,742/3,544 (77%) synthetic models that successfully convert, only 1,244/2,742 (45%) successfully load into ONNX Runtime. We observed 11 unique ONNX Runtime errors, of which 6 appear to correspond to open GitHub issues (vbogach, 2022; josephrocca, 2021; BowenBao, 2023; rafaelagrc, 2022; fatcat z, 2022; 12sf12, 2022). We disclosed the rest to the engineering teams. For most disclosures, we have not yet received a response. For one disclosure, which described a model that causes the libc++ abi to terminate unexpectedly during inference with ONNX Runtime, it had been (unintentionally) fixed in the development branch. The defect still affects older releases.
@@ 6.2.3. Method for Micro Analysis
For operator types, we measure the operators present in each converted model, out of the set of available operators in the ONNX opset. For operator sequences, we extract the simple paths through each model. To illustrate, the ONNX model in Figure 2 shows three operators (ReduceMax_0, ArgMax, and ReduceMax) and two simple linear paths (ReduceMax_0ArgMax and ReduceMax_0 ReduceMax). Models with similar architectures may have simple paths that are largely identical, inflating the number of shared sequences. For example, the sequences aaaabd and aaaabc are identical except for the last element. To deal with cases we further reduce the common sequences to the smallest shared subsequences, i.e., we recursively find the longest common subsequences for our sequences until we cannot find any smaller subsequences (Cormen et al., 2022). After reduction we find that sequences are between 3 to 5 operators long.
@@ 7.1. Validating DL Model Converters
Converter test suites share little in common with failing models, often missing critical model structures (section 6.2.4). Specifically, model converters need better testing techniques for behavioral changes. Though this critical failure mode occurred in 1/3 of historical failures (section 5.2) and affected 6% of the models we tested (section 6.2.2), the existing test suite models share little in common with models affected by this failure mode. For example, only 37 operator sequences are shared between mismatching models and test suite models (Table 12). Introducing operator sequence coverage metrics may help improve the quality of test suites, to promote the diversity of the models tested. These metrics could also be incorporated into fuzzers. Since some nodes or sequences are more likely to cause failures (Table 12), a coverage-guided fuzzing approach could help identify them.
Identifying behavioral changes after model conversion is critical, this is typically done with an end-to-end test. End-to-end testing carries with it the need to assess the outcome: is the converted model acceptable, and how much error is tolerable (Jiang et al., 2024a; Face, 2023)? Engineers currently use a variety of heuristic tolerances derived from experience (section 6.2.1). Theoretical and empirical examination of the expected and acceptable tolerance would benefit DL model converter testing.
@@ 7.2.1. DL Model Converters
Openja et al. evaluated DL model converters by converting 5 models (Openja et al., 2022). We extend their findings in two directions. First, we conducted a failure analysis (section 5) to study how converters fail. Second, we convert models systematically (section 6) and analyze them to understand why they fail. We considered both realistic and synthetic models. For realistic models, rather than picking 5 real-world models, we used 1,605 models from a large model corpus. For synthetic models, we adapted a DL model generation tool to conduct a bounded systematic exploration of converter behaviors. We omitted measurements of model size, adversarial robustness, and prediction accuracy, to focus instead on measuring the common failure modes (crashing and behavioral differences) identified in our failure analysis. Our analysis reveals that converters can successfully convert many real model but synthetic models are more prone to failure, moreover analyzing the synthetic models reveals that particular operator sequences co-occur with failures.
@@ 7.2.2. DL Compilers
Shen et al. studied failures in DL compilers (Shen et al., 2021), including the components that load models from DL frameworks and from common intermediary formats such as ONNX. Many of their results generalized to DL model converters: (1) we were able to adapt their taxonomies of failure symptoms and causes to DL model converters (section 5.1.1, Table 7); and (2) as shown in Table 9, we found a similar distribution of failure symptoms for DL model converters. However, the causes of failures between the two contexts differed: as noted in section 5.2.3, Incompatibilities were more common in DL model converters, while Algorithmic Errors were more common in DL compilers.
@@ 7.2.3. Methodological Innovations on the “Bug Study”
Our study began with a survey that offered valuable insights into the problem domain, guiding our design of the subsequent failure study, including the selection of the interoperability tool and the creation of a failure taxonomy (section 5). This methodological progression, from survey to failure study, and then to hypothesis testing about the failures, diverges from the typical failure study in software engineering. Those studies typically focus on the “middle step” (Amusuo et al., 2022), understanding the root causes of defects and the conditions under which they manifest (e.g., (Islam et al., 2019; Zhang et al., 2019)). Our hypothesis testing phase, in particular, adds understanding by providing an estimate of failure rates, an aspect seldom addressed in standard failure studies. We suggest that this holistic approach, integrating a survey, failure study, and hypothesis testing, may be a useful methodology for future failure studies.
@@ 8. Threats to Validity
We discuss three types of threats to validity (Wohlin et al., 2012). Taking into consideration the criticisms of Verdecchia et al. (Verdecchia et al., 2023), we focus on substantive threats that might influence our findings.
Construct Threats are potential limitations of how we operationalized concepts. We mitigated definitional concerns for RQ2 by adapting existing taxonomies in our failure analysis. We defined failures consistently with other works (GitHub issues closed with a repair), though we note there are other means of failure discovery (Aranda and Venolia, 2009). To answer RQ3, we conservatively defined failure in terms of clear misbehaviors — when ONNX converter output is incompatible with ONNX Runtime, or the model’s behavior changes. This may mask some “Crash” failures in the ONNX converter. In addition, different measures of behavioral difference are possible, such as the L1- or L2- distance. We used the measure recommended by PyTorch.
Internal threats are those that affect cause-effect relationships. We specifically allowed respondents to leave answers blank to improve the response rate. As a result, the results for interoperability tool use in RQ1 may be biased as only 48/92 respondents use an interoperability tool (section 4.1.2). However, of the 48 respondents who reported using interoperability tools, 98% (47/48) gave analyzable data.
Our failure characterization in RQ2 was manual. We mitigated subjectivity via interrater agreement, on a pilot sample and a subsequent tranche during the full analysis (section 5.1.2). In section 6.1.1 we noted several approximations in our test of hypothesis H. We tested hypothesis H indirectly, using a frequency analysis of operators and sequences in mismatched models rather than directly testing the discovered subsequences.
External threats may impact generalizability. For RQ1, the surveyed population (HuggingFace users) may not fully reflect all ONNX users. However, it is an appropriate population, and we sampled at a confidence level of 90%. For RQ2-4, we examined two DL model converters for one interoperability framework (ONNX). Our results may not extend to other model converters nor other interoperability tasks. As mitigation, our data suggest that ONNX is the most popular DL interoperability tool. For ONNX, our results were similar across the two converters, suggested generalizability in two dimensions: DNN modeling approaches, and converter owners (section 5.1.1). With regard to our methodology for RQ3-4, we generated synthetic models with the NNSmith tool. Other methods (Gu et al., 2022; Li et al., 2023; Wang et al., 2020) might have different results.
@@ 9. Conclusion
DL model converters play a crucial role in providing interoperability between DL frameworks, registries, and runtimes. Understanding the nature of failures in DL model converters enables engineers to make informed decisions when relying on them. We conducted the first failure analysis of DL model converters, considering the PyTorch and TensorFlow converters for the popular ONNX intermediate representation. The most common symptoms of failure are crashes and, perhaps more concerningly, models that misbehave on certain inputs. Of the five stages of a typical model converter, one stage (node conversion) accounts for 75% of the defects. A deductive description of the causes of erroneous converter behavior remains elusive — individual operators are not predictive of failure; sequences of operators may be correlated. Our findings suggest that in ONNX, engineers can rely on model converters but should validate the result for behavioral consistency. Through a mix of positive and negative results, we exposed several directions for further improvement of DL model converters. The main opportunities are new measurements: a new architectural coverage measure for a DL model converter, and a refined measure of tolerance after conversion.
10. Data-Availability Statement
This paper is accompanied by an artifact (jaj, 2024) to support replication and reproduction. The artifact includes the survey data for Theme 1, the failure analysis data for Theme 2, and the processed results of data associated with the hypothesis evaluation for Theme 3. The artifact also includes the modified NNSmith model generator, and the specific synthetic models used in the study. The source code for the artifact is available at https://github.com/PurdueDualityLab/issta24-onnx-artifact.
Acknowledgements
This work is supported by Cisco and Google, as well by NSF awards #2107020, #2107230, #1813935, and #2104319.